OpenStudy (anonymous):

Solve 4x2 = 12x + 40. i can't find the two numbers -____-

OpenStudy (anonymous):

Is that 4 times 2 or 4x times 2?

OpenStudy (anonymous):

4x^2 2 is the exponent of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it is a quadratic? 4x^2-12x+40=0

OpenStudy (anonymous):

4x² -12 x - 40 = 0 | divde by 4 x² - 3x - 10 = 0 x²-3x-10 = (x-5) * (x+2) Hence your roots are 5 and -2

OpenStudy (anonymous):

@Stiwan but it's a quadratic.and -5 and 2 added don't equal -40 and multiplied don't equal -12?

OpenStudy (anonymous):

@jess_in_tenesee yes

OpenStudy (anonymous):

try plugging it into the quadratic formula?

OpenStudy (anonymous):

Did you remember you factored out 4?

OpenStudy (anonymous):

oooooh! that's why! i forgot to factor 4. ok hold on let me try that

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

lol ok so my final answer is x=5 and x=-2 (:

hero (hero):

That's because x^2 - 3x - 10 was factored, not 4x^2 - 12x - 40. If you factor 4x^2 - 12x - 40, you get: 4x^2 - 12x - 40 4x^2 - 20x + 8x - 40 4x(x - 5) + 8(x - 5) (4x + 8)(x - 5) Notice, from this you get back to the original quadratic. -20 and 8 are your two numbers. But that simplifies further to 4(x + 2)(x - 5)

hero (hero):

Your two numbers need to multiply to get -160 and add to get -12 a + b = -12 ab = 4(-40)

OpenStudy (anonymous):

Thank you ! (: @Hero